At a track meet the longest shot put throw by a boy is 25 feet and 8 inches. The longest shot put by a girl is 19 feet and 3 inches. How many times greater is the longest shot put throw by a boy than a girl?

Answer by Maths68(1474)

Boy record = 25 feet 8 inches
Girl record = 19 feet 19 inches
Boy record / Girl record = 25 feet 8 inches / 19 feet 19 inches
Change the feet into inches
1 feet = 12 inches
25 feet = 25* 12 = 300 inches + 8 inches = 308 inches
19 feet = 19 * 12= 228 inches + 19 inches = 247 inches
Boy record / Girl record = 308 inches / 247 inches
Boy record / Girl record = 1.246 inches
Boy record = 1.246 Girl record
So
it is 1.246 times greater than girls record.